OpenStudy (asdfghj8):

Solve for w. (w+2)^2 = 2w^2 +2w-4

OpenStudy (asdfghj8):

\[(w+2)^{2}=2w ^{2}+2w-4\]

Vocaloid (vocaloid):

start by expanding (w+2)^2 what do you get?

OpenStudy (asdfghj8):

w^2+4x+4

Vocaloid (vocaloid):

good! now we write: w^2+4x+4 = 2w^2 +2w-4 then we combine like terms

Vocaloid (vocaloid):

w^2 - 2w - 8 = 0

OpenStudy (asdfghj8):

i understand that part then we factor it out after this right?

Vocaloid (vocaloid):

yes, good! do you know what that factors to?

OpenStudy (asdfghj8):

2 and -4?

Vocaloid (vocaloid):

good! so we have (w-4)(w+2) = 0 so w = ?

OpenStudy (asdfghj8):

w=-2 and 4?

Vocaloid (vocaloid):

good job!

OpenStudy (asdfghj8):

Thank you so so so much!!!!!